According yo its computer, a robot that left its closet and travelled 1200m, had an average speed of 20.0 m/s. How long did the trip take?


1
Expert's answer
2021-03-10T17:22:02-0500
"t=\dfrac{d}{v}=\dfrac{1200\ m}{20\ \dfrac{m}{s}}=60\ s."